OpenStudy (anonymous):

A rectangular box is 24 inches long, 10 inches wide, and 15 inches high. if exactly 60 smaller identical rectangular boxes can be stored perfectly in this larger box, which of the following could be the dimensions, in inches, of these smaller boxes a. 2 by 5 by 6 b. 3 by 4 by 6 c 3 by 5 by 6 d. 4 by 5 by 6 e. 5 by 6 by 12 @whpalmer4

OpenStudy (anonymous):

We are not allowed to give you the answer, but we can help you understand it.

OpenStudy (anonymous):

Yea i know

OpenStudy (anonymous):

okay so if the large box equals 3,600 then all you need to do is do bxhxl to each one then multiply it by 60, which ever one equals 3,600 after doing lxwxh then multiplied by 60 is your answer.

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

ya np, so what did you get for your answer?

OpenStudy (anonymous):

volume of box =24*10*15=3600 volume of each box=3600/60=60

OpenStudy (anonymous):

A

OpenStudy (anonymous):

you are correct.

OpenStudy (anonymous):

Ok, Thank you :D

OpenStudy (anonymous):

ya np